# NdLinear: Don’t Flatten! Building Superior Neural Architectures by Preserving N-D Structure

### v2

@@ Abstract
Many high-impact machine learning tasks involve multi-dimensional data such as images, volumetric medical scans, and multivariate time-series. Yet, most neural architectures flatten these inputs, discarding critical cross-dimension information. We introduce NdLinear, a novel linear transformation that circumvents this destructive flattening by operating directly on tensors. NdLinear applies transformations separately along each data dimension, thereby preserving the native data structure. Extensive experiments demonstrate NdLinear’s capacity to significantly enhance representational power, achieve dramatic parameter reductions (often by orders of magnitude), and maintain a favorable computational profile. For instance, when applied to Large Language Model finetuning, our NdLinear-LoRA delivers comparable or improved accuracy on reasoning tasks using up to 9×9$\times$9 × fewer trainable parameters than standard LoRA. These broad advantages of NdLinear are consistently validated across diverse neural architectures (CNNs, RNNs, Transformers, MLPs) and data domains, including vision, language, time-series, and tabular tasks. As a versatile, drop-in replacement for standard linear layers, NdLinear processes data in its original N-dimensional form, offering a foundational component for developing more efficient and powerful next-generation neural architectures.
@@ 1 Introduction
Effectively processing N-dimensional (N-D) tensors is critical in modern artificial intelligence, with applications spanning volumetric medical scans (Milletari et al., 2016), spatio-temporal video, multivariate time series (Liu et al., 2023), and outputs of complex simulations (Shanahan et al., 2023; Brown et al., 2020; Dosovitskiy et al., 2021; Radford et al., 2021). However, a core challenge remains: fully retaining and leveraging the rich, inherent structure of N-D data within neural architectures. While specialized layers like convolutions excel at capturing local spatial hierarchies (e.g., in 2D/3D images), they are not universally optimal, especially for comprehensively propagating information in higher-dimensional data (e.g., >3D) without excessive complexity. Fundamentally, many neural architectures (CNNs, Transformers, MLPs and RNNs) still rely on standard Linear (fully connected) layers, which expect 1D vector inputs. This common practice necessitates flattening multi-dimensional feature maps or tensors before these transformations, for instance, in network heads, channel-wise projections, or MLP blocks within attention mechanisms.
This ubiquitous flattening, even if localized, is often detrimental. It discards the data’s explicit N-D organization, disrupting critical spatial, temporal, or other relationships encoded across dimensions (Sokar et al., 2024). For example, valuable spatial hierarchies preserved by convolutions can be lost when feature vectors are decontextualized by flattening. In multivariate time series (e.g., dimensions for time, sensors, features), flattening intermingles these distinct axes, forcing models to rediscover temporal patterns or inter-sensor dependencies without structural priors. This loss of explicit structure forces networks to relearn intrinsic dependencies, potentially increasing parameter counts, data requirements, and computational load, while risking spurious correlations.
To address this, we introduce NdLinear, a novel linear layer that operates natively on N-D tensors. Unlike standard linear layers, NdLinear processes data in its original N-D form by applying distinct linear transformations sequentially along each tensor dimension. This dimension-wise approach inherently preserves the data’s N-D structure.
Preserving data’s native organization allows NdLinear to offer compelling advantages: It enhances representational power by maintaining structural integrity and facilitating a more natural information flow. It dramatically reduces parameter counts (often by orders of magnitude vs. flattened layers) and ensures efficient computation without sacrificing performance. These efficiencies lead to faster training/inference and lower memory usage. As a general-purpose layer, NdLinear is a versatile, drop-in replacement for standard linear layers in many architectures, benefiting diverse domains by enabling models to better leverage N-D data characteristics.
@@
  • We empirically demonstrate through comprehensive evaluations that NdLinear significantly enhances model performance, or matches it with drastically fewer parameters, when replacing standard linear layers across diverse architectures (CNNs, RNNs, Transformers, MLPs) and data domains, including vision, language, time-series, and tabular data.
@@
  • We establish NdLinear as a versatile and foundational building block for developing more efficient, structurally-informed, and powerful neural architectures.
The remainder of this paper is organized as follows. Section 2 details the NdLinear transformation. Section 3 analyzes its parameter efficiency, computational complexity, and representational capacity. Section 4 presents experimental results demonstrating NdLinear’s efficacy. Finally, Section 5 summarizes findings, discusses implications, and outlines future work.
2 NdLinear: N-dimensional Linear Transformation
This section formally introduces the NdLinear transformation. We first define the general problem of transforming tensor-structured data. Then, we detail the NdLinear layer, including its operational mechanics and algorithm. Finally, we situate NdLinear within the context of related methods.
2.1 Problem Formulation
This work addresses learnable linear transformations for N-dimensional (N-D) input tensors. Such an input, $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$, comprises a batch of $B$ samples, each with $N$ feature dimensions of sizes $D_{1},\ldots,D_{N}$ (e.g., height, width, channels for images; sequence length, embedding dimensions for text). The objective is to transform $X$ into an output tensor $Y^{\prime}\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$, where each $k$-th feature dimension is reshaped from its original size $D_{k}$ to a new size $H_{k}$. Our focus is specifically on the formulation of this N-D linear transformation.
A primary challenge is that conventional linear layers necessitate flattening $X$ into a 2D matrix (batch size $\times$ total features). This operation discards the tensor’s explicit N-D structure. We aim to define a linear transformation that operates directly on N-D tensors, thereby preserving their inherent multi-dimensional organization.
2.2 The NdLinear Transformation
The NdLinear transformation processes N-D tensors by applying separate linear transformations along each of their feature dimensions sequentially. This contrasts with standard linear layers, which flatten the N-D tensor into a 2D matrix, losing the original multi-dimensional structure. By operating on each dimension independently while preserving the others, NdLinear retains the inherent structural information of the data throughout the transformation.
@@ 2.2 The NdLinear Transformation
This sequential application can be expressed using tensor notation as a series of mode-$k$ tensor-matrix products (Kolda and Bader, 2009): $Y=X\times_{1}W_{1}\times_{2}W_{2}\cdots\times_{N}W_{N}$ , where each product $X\times_{k}W_{k}$ transforms the $k$-th mode of the tensor using matrix $W_{k}$. (Biases $b_{k}$ are added after each mode-$k$ product). The intermediate result of $X\times_{k}W_{k}$ becomes the input for the ×$\times_{k+1}W_{k+1}$ product.
2.3 Relationship to Existing Methods
NdLinear’s design as a structure-preserving, parameter-efficient linear layer for N-D tensors is informed by, yet distinct from, several established concepts in machine learning and tensor algebra.
First, NdLinear’s core mechanism, which is sequential matrix multiplications along each tensor mode, employs fundamental operations from multilinear algebra, prominent in tensor decompositions like Tucker decomposition (Tucker, 1966; Kolda and Bader, 2009). However, unlike classical tensor decomposition methods that primarily analyze or compress static data or pre-existing weight tensors (e.g., (Novikov et al., 2015), (Newman et al., 2024)), NdLinear integrates these mode-wise operations as a learnable, dynamic layer within a network. Its purpose is the efficient forward transformation of activations while preserving N-D structure, not data analysis or post-hoc model compression.
Second, the principle of factorizing operations along data axes for efficiency and structured inductive bias is shared with successful neural network components like depthwise separable convolutions (Chollet, 2017b; Howard et al., 2017b) (separating spatial/channel operations) and axial attention mechanisms (Ho et al., 2019; Wang et al., 2020; Yan et al., 2023) (processing sequence dimensions independently). NdLinear generalizes this factorization: it applies to any N-D tensor and factorizes the linear transformation along all of its modes, not being confined to specific mechanisms or predefined splits (e.g., spatial vs. channel).
Finally, NdLinear belongs to the family of structured linear layers designed for parameter efficiency (Denil et al., 2013), (Wei et al., 2024). While many approaches impose specific algebraic structures on a large conceptual weight matrix (e.g., low-rank (Sainath et al., 2013), or Monarch (Dao et al., 2022) and other forms (Sindhwani et al., 2015), (Potapczynski et al., 2024)), NdLinear’s structure is directly and interpretably derived from the N-D organization of the input data. Factorization occurs along the data’s native modes, offering a data-centric approach to efficient N-D linear transformation.
In essence, NdLinear synthesizes these inspirations into a versatile layer that directly applies linear transformations to N-D data without destructive flattening, prioritizing structural integrity and computational efficiency.
Comparison to Tensor Contraction and Tensor Regression Layers. Tensor Contraction Layers (TCLs) (Cichocki, 2014; Novikov et al., 2015) and Tensor Regression Layers (TRLs) (Kossaifi et al., 2020) also leverage tensor operations. TCLs often focus on dimensionality reduction by contracting specific modes. TRLs extend regression to multi-dimensional data using tensor formats. In contrast, NdLinear aims to transform N-D tensors while fully maintaining their N-D structure throughout the process, rather than primarily for dimensionality reduction or regression. NdLinear leverages the data’s intrinsic N-D nature for parameter efficiency, offering a complementary approach.
3 Analysis of NdLinear: Efficiency and Representational Properties
NdLinear’s design (Section 2) yields significant efficiency advantages and beneficial inductive biases for N-D data. Here, we quantify its parameter and computational efficiency against standard linear layers and examine how its structure shapes representational capacity and guides learning.
3.1 Drastic Parameter Reduction
A hallmark of NdLinear is its exceptional parameter economy. Consider transforming an input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ to an output $Y^{\prime}\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$. A conventional linear layer operating on flattened tensors (input size $D_{\text{flat}}=\prod D_{k}$, output size $H_{\text{flat}}=\prod H_{k}$) requires $D_{\text{flat}}\cdot H_{\text{flat}}+H_{\text{flat}}$ parameters (weights and biases).
NdLinear circumvents this combinatorial growth. By employing $N$ mode-specific transformations, each with a weight matrix $W_{k}\in\mathbb{R}^{D_{k}\times H_{k}}$ and optional bias $b_{k}\in\mathbb{R}^{H_{k}}$, NdLinear’s total parameter count is only $\sum_{k=1}^{N}(D_{k}H_{k}+H_{k})$.
The practical impact is profound. For example, transforming a 3D tensor where all $D_{k}=H_{k}=D=32$, a standard linear layer requires $O(D^{2N})$ parameters, which is $32^{6}\approx 10^{9}$ parameters. In contrast, NdLinear requires $O(ND^{2})$ parameters, merely $3\cdot 32^{2}\approx 3\times 10^{3}$, a reduction of several orders of magnitude. This drastic reduction makes NdLinear highly advantageous for leaner models, reduced memory usage, and potentially mitigating overfitting, especially with high-dimensional N-D inputs.
3.2 Efficient Computation
Corresponding to its parameter economy, NdLinear demonstrates significant computational efficiency. A standard linear layer performs approximately $2B\cdot(\prod D_{k})\cdot(\prod H_{k})$ Floating Point Operations (FLOPs).
NdLinear involves $N$ sequential mode-wise transformations. The total FLOPs are:
$\text{FLOPs}_{\text{NdLinear}}\approx 2B\sum_{k=1}^{N}\left[\left(\prod_{j=1}^{k-1}H_{j}\right)\left(\prod_{j=k+1}^{N}D_{j}\right)D_{k}H_{k}\right].$ (1)
This sum is typically far smaller than the product for a standard layer. Using the previous example ($N=3,D_{k}=H_{k}=D=32$), a standard layer requires $O(BD^{2N})$ FLOPs ($\approx B\cdot 10^{9}$). NdLinear requires $O(NBD^{N+1})$ FLOPs ($\approx B\cdot 3\cdot 10^{6}$). This represents orders of magnitude in computational savings, making NdLinear significantly faster, especially as $N$ or dimension sizes $D_{k},H_{k}$ increase. Its sequential operations also allow for efficient memory management of intermediate results.
3.3 Structured Representation and Beneficial Inductive Bias
NdLinear’s efficiency stems directly from its structured, factorized transformation of N-D tensors, a design that inherently shapes its representational capabilities and introduces inductive biases.
A single NdLinear layer, with its mode-wise factorization and reduced parameter set, represents a constrained subset of all possible linear transformations. This constraint is its primary inductive bias: NdLinear is predisposed to learn "separable patterns" aligned with the data’s intrinsic axes through sequential, dimension-wise transformations. This bias is particularly potent for data with inherent modal structures (e.g., temporal, spatial, channel-wise), guiding the model to leverage this organization for more efficient learning and better generalization than possible from flattened vectors.
This structural constraint also induces strong regularization. By limiting model complexity compared to dense layers, NdLinear naturally mitigates overfitting, especially with limited training data. This is akin to benefits from explicit complexity reduction (e.g., low-rank assumptions (Sainath et al., 2013; Denil et al., 2013)), achieved here through NdLinear’s intrinsic low-complexity design.

### v3

# NdLinear: Preserving Multi-Dimensional Structure for Parameter-Efficient Neural Networks

@@ Abstract
In deep learning, processing multidimensional inputs (e.g., images, medical scans, and time series) is an important task that often requires flattening the inputs. We introduce NdLinear, a drop-in replacement for linear layers that operates directly on tensors, requiring no flattening. By applying transformations separately along each dimension, NdLinear preserves native data structure while achieving dramatic parameter reductions, often by orders of magnitude, with minimal memory overhead. We prove NdLinear maintains expressivity through structured Tucker decomposition while preserving VC-dimension scaling. Extensive experiments demonstrate NdLinear’s capacity to achieve significant parameter reductions with substantial wall-clock efficiency gains and minimal memory overhead. For instance, our NdLinear-LoRA matches or exceeds standard LoRA on language reasoning tasks using up to $9\times$ fewer parameters. Experiments across CNNs, RNNs, Transformers, and MLPs on vision, language, time-series, and tabular tasks consistently demonstrate NdLinear’s efficiency gains. While excelling at axis-separable tasks, NdLinear has limitations with entangled spatial interactions. By processing data in its original N-dimensional form, NdLinear provides a theoretically grounded, practical component for building more efficient neural architectures.
@@ 1 Introduction
Deep learning excels at processing multidimensional data, such as medical scans, videos and sensor arrays. Yet, a fundamental inefficiency persists throughout modern architectures. Consider what happens when a 32×32×32 voxel tensor reaches a linear layer: it gets flattened into a 32,768-dimensional vector, scrambling the careful spatial organization that earlier layers worked to extract. This pattern repeats everywhere linear layers appear: CNN classification heads flatten spatial feature maps, Transformer MLPs flatten structured token representations, and RNN projections flatten temporal features. Beyond this massive inefficiency, flattening forces networks to waste capacity rediscovering that adjacent voxels are related, that temporal continuity matters, and that feature dimensions have distinct meanings. While convolutions preserve local structure, they cannot replace linear layers for global reasoning, channel mixing, or final predictions. This leaves a critical gap: no existing layer can perform general linear transformations while respecting multidimensional structure.
To address this, we introduce NdLinear, a novel linear layer that operates natively on N-D tensors. Unlike vanilla linear layers, NdLinear processes data in its original N-D form by applying distinct linear transformations sequentially along each tensor dimension. This dimension-wise approach inherently preserves the data’s N-D structure.
Preserving data’s native organization allows NdLinear to offer compelling advantages: It enhances representational power by maintaining structural integrity and facilitating a more natural information flow. It dramatically reduces parameter counts (often by orders of magnitude vs. flattened layers) and ensures efficient computation without sacrificing performance. These efficiencies lead to faster training/inference and lower memory usage. As a general-purpose layer, NdLinear is a versatile, drop-in replacement for vanilla linear layers in many architectures, benefiting diverse domains by enabling models to better leverage N-D data characteristics.
@@
  • We empirically demonstrate through comprehensive evaluations that NdLinear significantly enhances model performance, or matches it with drastically fewer parameters, when replacing vanilla linear layers across diverse architectures (CNNs, RNNs, Transformers, MLPs) and data domains, including vision, language, time-series, and tabular data.
@@
By demonstrating that preserving N-D structure yields both theoretical and practical advantages, NdLinear challenges the conventional wisdom that flattening is a necessary evil in neural architectures.
2 Related Work
Neural networks process multidimensional data through three main approaches, each with fundamental limitations:
Flattening-based Methods.
Standard linear layers reshape N-D tensors to 2D matrices, destroying spatial and dimensional relationships. This forces networks to relearn structure from scrambled features, requiring excessive parameters. For example, a $32\times 32\times 32$ tensor needs $\sim 10^{9}$ parameters when flattened, versus $\sim 10^{3}$ if structure were preserved.
Specialized Architectures.
Convolution layers excel at 2D/3D spatial patterns but become unwieldy beyond three dimensions. Depthwise separable convolutions (Chollet, 2017a) and axial attention (Ho et al., 2019) reduce complexity through factorization but remain tied to specific spatial operations rather than general tensor transformations.
Tensor Decomposition Layers.
Recent work applies tensor algebra to neural networks. Tensor Contraction Layers (TCL) (Kossaifi et al., 2020) use multilinear maps to compress activations to lower dimensions. Tensor Regression Layers (TRL) (Kossaifi et al., 2020) parameterize predictions using Tucker or Tensor-Train formats. Tensor-Train layers (Novikov et al., 2015) decompose weights for compression. While efficient, these methods focus on specific tasks, dimensionality reduction or regression, rather than general-purpose transformation.
Gap in Current Methods.
Despite extensive work on structured layers, no method provides a true drop-in replacement for linear layers that: (1) operates directly on N-D tensors without flattening, (2) supports flexible dimension-wise transformation (expansion or compression), (3) integrates seamlessly with modern architectures (bias, normalization, dropout). This gap motivates NdLinear, which we introduce in Section 3.
We provide detailed mathematical comparisons between NdLinear and existing tensor methods (TCL, TRL, $T$ decomposition) in Appendix A, demonstrating how our approach differs fundamentally in design philosophy and implementation, leading to the superior empirical results shown in Section 4.
3 Linear Transformation Preserving N-dimensional Information
Vanilla linear layers cannot process input tensors directly. They require transforming the inputs into 2D matrices, destroying the original N-D structure. For an N-D input tensor $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ with batch size $B$ and feature dimensions $(D_{1},\ldots,D_{N})$, standard layers flatten it to $\mathbb{R}^{B\times\prod_{i}D_{i}}$ before applying a linear transformation. NdLinear suggests an alternative approach that transforms $X$ directly to $Y\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$ without flattening, where each dimension $D_{k}$ maps to $H_{k}$, while preserving the N-D structure throughout the transformation.
The NdLinear transformation processes N-D tensors by applying separate linear transformations along each of their feature dimensions sequentially. This contrasts with vanilla linear layers, which flatten the N-D tensor into a 2D matrix, losing the original multidimensional structure. By operating on each dimension independently while preserving the others, NdLinear retains the inherent structural information of the data throughout the transformation.
@@ 3 Linear Transformation Preserving N-dimensional Information
This sequential application can be expressed using tensor notation as a series of mode-$k$ tensor-matrix products (Kolda and Bader, 2009): $Y=X\times_{1}W_{1}\times_{2}W_{2}\cdots\times_{N}W_{N}$ , where each product $X\times_{k}W_{k}$ transforms the $k$-th mode of the tensor using matrix $W_{k}$. (Biases $b_{k}$ are added after each mode-$k$ product). The intermediate result of $X\times_{k}W_{k}$ becomes the input for the $\times_{k+1}W_{k+1}$ product.
3.1 Preserving The Expressiveness of Vanilla Linear Layers
The dramatic parameter reduction of NdLinear raises a natural question: does this efficiency sacrifice model expressivity? We show that despite using fewer parameters, NdLinear maintains sufficient representational capacity.
Theoretical Guarantee. We analyze expressivity through VC-dimension, which measures a model’s capacity to fit arbitrary patterns. Following Bartlett et al. (2019), networks with $P$ parameters have VC-dimension $\Theta(P\log P)$.
Theorem 3.1 (Informal; see Appendix C for formal statement).
An NdLinear network with $P_{\text{nd}}=d(a+b+c)$ parameters for tensor dimensions $(a,b,c)$ and hidden dimension $d$ maintains VC-dimension $\Theta(P_{\text{nd}}\log P_{\text{nd}})$ as $d\to\infty$, matching the scaling of vanilla linear layers with $P_{\text{std}}$ parameters.
Practical Implication. While NdLinear uses fewer parameters, the reduction is polynomial and not exponential in the VC-dimension bound. This theoretical guarantee is validated empirically: NdLinear often improves performance despite parameter reduction, suggesting the structured factorization acts as beneficial regularization rather than a limiting constraint.
Empirical Evidence: Representation Compression. To understand why fewer parameters improve performance, we measured von Neumann entropy of learned representations. On the Radius Bump task F.3 across varying difficulties, NdLinear consistently produces 15-30% lower entropy than parameter-matched dense networks while achieving equal or better test MSE (Appendix F.3). This lower entropy, indicating more compressed representations, aligns with recent theoretical work showing compressed representations generalize better (Skean et al., 2025). NdLinear’s structured factorization inherently eliminates redundancy while preserving task-relevant information, explaining why dramatic parameter reduction enhances rather than hurts performance.
3.2 Fewer Learnable Parameters And Wall-Clock Speedups
A key advantage of NdLinear is its parameter efficiency. Consider transforming $X\in\mathbb{R}^{B\times D_{1}\times\cdots\times D_{N}}$ to $Y\in\mathbb{R}^{B\times H_{1}\times\cdots\times H_{N}}$.
Parameter Reduction. Vanilla linear layers require $(\prod_{i}D_{i})\times(\prod_{i}H_{i})$ parameters, which grow exponentially with dimensionality. NdLinear requires only $\sum_{i=1}^{N}(D_{i}H_{i})$ parameters, which grow linearly. For example, transforming a $32\times 32\times 32$ tensor to the same size requires $\sim 10^{9}$ parameters for vanilla linear but only $3,072$ for NdLinear, a reduction of six orders of magnitude.
Computational Complexity. NdLinear’s FLOPs scale as $\mathcal{O}(B\cdot N\cdot D^{N+1})$ for cubic tensors with $D_{i}=H_{i}=D$, compared to $\mathcal{O}(B\cdot D^{2N})$ for vanilla linear layers. This yields order-of-magnitude speedups that increase with tensor dimensionality. The exact FLOP count is:
$\text{FLOPs}_{\text{NdLinear}}=B\sum_{k=1}^{N}\left(\prod_{j<k}H_{j}\cdot\prod_{j>k}D_{j}\cdot D_{k}H_{k}\right).$
Memory Efficiency. While theoretical analysis bounds peak memory overhead at $1/N$ of baseline, which is typically $\leq 33\%$ for 3D tensors (proof in Appendix C.2), empirical measurements show much smaller overhead in practice: only 1.1-2.0% increased peak memory across diverse architectures, with training time overhead below 1.6% (Section 4.5).
3.3 Inductive Bias and Domain Alignment
NdLinear’s efficiency stems from a strong inductive bias: it assumes dimensions can be transformed independently. This bias has clear benefits and limitations.
The Separability Assumption. NdLinear implements a rank-1 Tucker decomposition, assuming the transformation can be factorized across dimensions. Crucially, this structure persists through ReLU-like activations. For 1-homogeneous activations $\sigma$ (ReLU, GELU):
$\sigma(T(x))=\sum_{r=1}^{R}\bigotimes_{k=1}^{m}\sigma(u_{r}^{(k)}),$
where output rank remains bounded even after nonlinearities. This ensures the separability bias persists through network depth rather than degrading.
Domain Alignment. NdLinear excels when data dimensions represent independent factors, such as spectrograms (frequency × time), sensor arrays (sensor × time), and tabular data, where features have separable effects. However, it struggles with dense cross-dimensional interactions like XOR patterns, checkerboards, or highly entangled spatial features where dimensions cannot be meaningfully separated.
3.4 Quantifying the Trade-off: Dial-a-Bias Experiment
To precisely characterize when NdLinear’s bias helps versus hurts, we designed a controlled experiment interpolating between separable and entangled patterns.
We create synthetic targets that blend separable and entangled components:
$y=(1-\alpha)\cdot f_{\text{separable}}(X)+\alpha\cdot f_{\text{entangled}}(X),\quad\alpha\in[0,1],$
where $X\in[0,1]^{32\times 32}$ has i.i.d. uniform entries. The separable component $f_{\text{separable}}$ aggregates per-axis statistics (row/column means, variances), while $f_{\text{entangled}}$ requires cross-axis interactions (XOR patterns, checkerboards).
Figure 1: NdLinear excels on separable tasks but struggles with entangled patterns Performance comparison as task structure varies from purely separable ($\alpha=0$) to fully entangled ($\alpha=1$). The crossover at $\alpha\approx 0.45$ indicates NdLinear outperforms dense MLPs when tasks have <<45% entanglement. This provides clear deployment guidance: use NdLinear for axis-aligned domains (spectrograms, time series, tabular data) but prefer dense layers for spatially entangled tasks (dense vision, XOR-like patterns).
Experimental Protocol. We trained both NdLinear and dense MLPs (2 hidden layers, 256 units) on 10,000 samples for each $\alpha\in\{0,0.1,...,1.0\}$.
Results. Figure 1 reveals a sharp crossover at $\alpha\approx 0.45$. NdLinear achieves 49× lower MSE for purely separable tasks ($\alpha=0$) but 64× worse for fully entangled tasks ($\alpha=1$). This provides clear guidance on the use of NdLinear, where the target structure aligns with axis-separable patterns.